An ideal transformer has 4500. turns in its primary (input) and 14500. turns in its secondary (output). 10.0-V ac is connected to the primary (input).

a) What is the voltage on the secondary (output)?

b) What is the secondary (output) current if an 8.00-Ω resistor is connected to the secondary (output)?

c) What is the power supplied to the primary (input)?
Answer #1

a)

Np = number of turns in primary= 4500

Ns = number of turns in secondary = 14500

Vp = primary Ac Voltage = 10 volts

Vs = secondary voltage

Using the equation

Vs /Vp = Ns /Np

Vs /10 = 14500/4500

Vs = 32.2 Volts

b)

Rs = resistance connected to secondary side = 8 ohm

is = current on secondary side

Using ohm's law

Vs = is Rs

32.2 = is (8)

is = 4.025 A

c)

ip = current in primary side

Using the equation

ip/is = Vs/Vp

ip/4.025 = 32.2/10

ip = 12.96 A

P = Power in primary side = ip Vp = 12.96 x 10 = 129.6 Watt